OpenStudy (anonymous):

How do I solve this problem when a and b are assumed to represent any real number? ^418sqrt(a+b)^418

OpenStudy (anonymous):

ah what is \[\sqrt[2]{x ^{2}}\]

OpenStudy (anonymous):

x^4? lol I'm not sure

jimthompson5910 (jim_thompson5910):

If n is even and x is a real number, then \[\sqrt[n]{x^n}=|x|\] In your case, n=418 and x=a+b, so \[\sqrt[418]{(a+b)^{418}}=|a+b|\]

OpenStudy (anonymous):

because \[\sqrt{4} = \sqrt{2^{2}} = 2\]

OpenStudy (anonymous):

so the 418 is cancelled out because you are multiplying (a+b) 418 times and then taking the 418th root of said number, so the answer is the radicand.